### v1

# R-enum Revisited: Speedup and Extension for Context-Sensitive Repeats and Net Frequencies

Abstract
A repeat is a substring that occurs at least twice in a string, and is called a maximal repeat if it cannot be extended outwards without reducing its frequency. Nishimoto and Tabei [CPM, 2021] proposed r-enum, an algorithm to enumerate various characteristic substrings, including maximal repeats, in a string $T$ of length $n$ in $O(r)$ words of compressed working space, where $r\leq n$ is the number of runs in the Burrows-Wheeler transform (BWT) of $T$. Given the run-length encoded BWT (RLBWT) of $T$, r-enum runs in $O(n\log\log_{w}(n/r))$ time in addition to the time linear to the number of output strings, where $w=\Theta(\log n)$ is the word size. In this paper, we improve the $O(n\log\log_{w}(n/r))$ term to $O(n)$. We also extend r-enum to compute other context-sensitive repeats such as near-supermaximal repeats (NSMRs) and supermaximal repeats, as well as the context-diversity for every maximal repeat in the same complexities. Furthermore, we study the occurrences that witness NSMRs, which have recently attracted attention under the name of net occurrences: An occurrence of a repeat is called a net occurrence if it is not covered by another repeat, and the net frequency of a repeat is the number of its net occurrences. With this terminology, an NSMR is defined to be a repeat with a positive net frequency. Given the RLBWT of $T$, we show how to compute the set $\mathcal{S}^{\mathsf{nsmr}}$ of all NSMRs in $T$ together with their net frequency/occurrences in $O(n)$ time and $O(r)$ space. We also show that an $O(r)$-space data structure can be built from the RLBWT to support queries of computing the net frequency of any query pattern $P$ in optimal $O(|P|)$ time. The data structure is built in $O(r)$ space and in $O(n)$ time with high probability or deterministic $O(n+|\mathcal{S}^{\mathsf{nsmr}}|\log\log\min(\sigma,|\mathcal{S}^{\mathsf{nsmr}}|))$ time, where $\sigma\leq r$ is the alphabet size of $T$. To achieve this, we prove that the total number of net occurrences is less than $2r$. With the duality between net occurrences and minimal unique substrings (MUSs), we get a new upper bound $2r$ of the number of MUSs in $T$, which may be of independent interest.
1 Introduction
Finding characteristic substring patterns in a given string $T$ is a fundamental task for string processing. One of the simplest patterns based on frequencies would be a repeat, which occurs at least twice in $T$. Since the number of distinct repeats in $T$ of length $n$ can be $\Theta(n^{2})$, maximality is often utilized to reduce the number down to $O(n)$ [13]. A substring of $T$ is called right-maximal (resp. left-maximal) if it cannot be extended to the right (resp. to the left) without reducing its number of occurrences in $T$. If a repeat is both right-maximal and left-maximal, it is called a maximal repeat (MR).
Gusfield [13] also introduced stricter criteria to reduce the redundancy in maximal repeats. A repeat is called a supermaximal repeat (SMR) if it is not a substring of another repeat, and a near-supermaximal repeat (NSMR) if at least one of its occurrences is not covered by another repeat. It is easy to see that an SMR is an NSMR, and an NSMR is an MR. NSMRs were studied under the name of Chinese frequent strings in [17, 18, 28], and largest maximal repeats in [23, 7]. Gallé and Tealdi [8, 9] proposed the concept of context diversities to define a general class of context-sensitive repeats including MRs and SMRs.
Enumerating these context-sensitive repeats has numerous applications in bioinformatics [13, 23, 1] and natural language processing [17, 18, 29, 20, 7] especially valuable for languages with no clear word segmentation such as Chinese and Japanese. Motivated by these applications, efficient enumeration algorithms have been extensively studied [29, 1, 27, 2, 26] using a family of suffix data structures such as suffix trees [31], directed acyclic word graphs (DAWGs) [5], compact DAWGs (CDAWGs) [4, 16], suffix arrays [19] and Burrows-Wheeler transform (BWT) [6].
In this paper, we focus on r-enum [26], an algorithm to enumerate the MRs, minimal unique substrings (MUSs) [14] and minimal absent words (MAWs) in $O(r)$ words of compressed working space, where $r\leq n$ is the number of runs in the BWT of $T$ of length $n$. Given the run-length encoded BWT (RLBWT) of $T$, r-enum runs in $O(n\log\log_{w}(n/r))$ time in addition to the time linear to the number of output strings, where $w=\Theta(\log n)$ is the word size. We show that $O(n\log\log_{w}(n/r))$ time can be improved to $O(n)$ by resorting to the move data structure recently proposed by Nishimoto and Tabei [24] for constant-time LF-mapping on RLBWTs. We also extend r-enum to compute the NSMRs and SMRs, and the context diversity for every RMR in the same time and space complexities.
Furthermore, we consider problems for net occurrences and net frequencies, originated from the studies of Chinese frequent strings [17, 18] and have recently attracted attention from combinatorial and algorithmic points of view [10, 28, 12, 15, 11, 21]. An occurrence of a repeat is called a net occurrence if the occurrence is not covered by another repeat, and the net frequency of a repeat is the number of its net occurrences. We would like to point out the fact, which may have been overlooked for over two decades, that the strings with positive net frequencies are equivalent to the NSMRs [13] and the largest maximal repeats [23, 7], and thus appear to have broader applications. Also, net occurrences are conceptually equivalent to maximum repeats in [14].
Given the RLBWT of $T$, we show how to compute the set $\mathcal{S}^{\mathsf{nsmr}}$ of all NSMRs in $T$ together with their net frequency/occurrences in $O(n)$ time and $O(r)$ words of space, solving the task called All-NF in [10] on RLBWTs. We also show that an $O(r)$-space data structure can be built from the RLBWT to support queries of computing the net frequency of any query pattern $P$ in optimal $O(|P|)$ time. The data structure is built in $O(r)$ space and in $O(n)$ time with high probability or deterministic $O(n+|\mathcal{S}^{\mathsf{nsmr}}|\log\log\min(\sigma,|\mathcal{S}^{\mathsf{nsmr}}|))$ time, where $\sigma\leq r$ is the alphabet size of $T$. To achieve this, we prove that the total number of net occurrences is less than $2r$. With the duality between net occurrences and MUSs shown in [14, 21], we immediately get a new upper bound $2r$ of the number of MUSs in $T$, which may be of independent interest.
2 Preliminaries
2.1 Basic notation
We assume a standard word-RAM model with word size $w=\Theta(\log n)$, where $n$ is the length of string $T$ defined later. A space complexity is evaluated by the number of words unless otherwise noted. Any integer treated in this paper is represented in $O(1)$ space, i.e., $O(w)$ bits. A set of consecutive integers is called an interval, which can be represented in $O(1)$ space by storing its beginning and ending integers. For two integers $b$ and $e$, let $[b..e]$ denote the interval $\{b,b+1,\dots,e\}$ beginning at $b$ and ending at $e$ if $b\leq e$, and otherwise, the empty set. We also use [b..e):=[b..e1][b..e):=[b..e-1] to exclude the right-end integer.
A string (or an array) $x$ over a set $S$ is a sequence $x[1]x[2]\cdots x[|x|]$, where $|x|$ denotes the length of $x$ and $x[i]\in S$ is the $i$-th element of $x$ for any $i\in[1..|x|]$. The string of length 0 is called the empty string and denoted by $\varepsilon$. Let $S^{*}$ denote the set of strings over $S$, and let $S^{n}\subset S^{*}$ be the set of strings of length $n\geq 0$. For integers $1\leq b\leq e\leq|x|$, the substring of $x$ beginning at $b$ and ending at $e$ is denoted by $x[b..e]=x[b]x[b+1]\cdots x[e]$. For convenience, let $x[b..e]$ with $b>e$ represent the empty string. We also use x[b..e)=x[b..e1]x[b..e)=x[b..e-1] to exclude the right-end character. A prefix (resp. suffix) of $x$ is a substring that matches $x[1..e]$ (resp. $x[b..|x|]$) for some $e\in[0..|x|]$ (resp. $b\in[1..|x|+1]$). We use the abbreviation $x[..e]=x[1..e]$ and $x[b..]=x[b..|x|]$. For a non-negative integer $d$, let $x^{d}$ denote the string of length $|x|d$ such that $x^{d}[|x|(i-1)+1..i|x|]=x$ for any $i\in[1..d]$.
2.2 Tools
A function that injectively maps a set $S$ of integers to $[1..O(|S|)]$ can be used as a compact dictionary for $S$ to look up a value associated with an integer in $S$. We use the following known result:
Lemma 1 ([32, 30]).
Given $S\subseteq[1..u]$ of size $m\leq u\leq 2^{O(w)}$, we can build an $O(m)$-size dictionary in $O(m)$ time with high probability or deterministic $O(m\log\log m)$ time so that lookup queries can be supported in $O(1)$ worst-case time.
A predecessor query for an integer $i$ over the set $S$ of integers asks to compute $\max\{j\in S\mid j\leq i\}$, which can be efficiently supported by the data structure of [3, Appendix A]:
Lemma 2 ([3]).
Given $S\subseteq[1..u]$ of size $m\leq u\leq 2^{O(w)}$, we can build an $O(m)$-size data structure in $O(m\log\log_{w}(u/m))$ time and $O(m)$ space to support predecessor queries in $O(\log\log_{w}(u/m))$ time for any $i\in[1..n]$.
A range distinct query $\mathsf{RD}(x,p,q)$ for a string $x$ and integers $1\leq p\leq q\leq|x|$ asks to enumerate $(c,p_{c},q_{c})$ such that $p_{c}$ and respectively $q_{c}$ are the smallest and largest positions in $[b..e]$ with $x[p_{c}]=x[q_{c}]=c$ for every distinct character $c$ that occurs in $x[p..q]$. It is known that a range distinct query can be supported in output-optimal time after preprocessing $x$.
Lemma 3 ([22, 2]).
Given a string $x\in[1..\sigma]^{m}$, we can build an $O(m)$-size data structure in $O(m)$ time and space to support range distinct queries. With a null-initialized array of size $\sigma$, each query can be answered in $O(k)$ time, where $k$ is the number of outputs.
A minor remark is that the output of range distinct queries of Lemma˜3 may not be sorted in any order of characters.
2.3 Context-sensitive repeats
[…]
To describe algorithms uniformly, we assume that the empty string is a repeat that occurs at every position $b$ in $[1..n]$, i.e., $\mathsf{Occ}(\varepsilon)=[1..n]$. Since $\mathsf{lc}(\varepsilon)=\mathsf{rc}(\varepsilon)=\Sigma_{\mathdollar}$, $\varepsilon$ is always in $\mathcal{S}^{\mathsf{lmr}}$, $\mathcal{S}^{\mathsf{rmr}}$ and $\mathcal{S}^{\mathsf{mr}}$. We set the following exception for the definition of net occurrences of $\varepsilon$: An occurrence $b\in\mathsf{Occ}(\varepsilon)$ is a net occurrence of $\varepsilon$ if and only if T[b1]T[b-1] and $T[b]$ are both unique. This exception provides a smooth transition to another characterization of net occurrences based on the uniqueness of extended net occurrences [10, 21].
Table 2 summarizes the concepts introduced in this subsection for $T=\mathtt{abcbbcbcabc\mathdollar}$.
$x\in\mathcal{S}^{\mathsf{r}}$ occurrences $\mathsf{Occ}(x)$ $\mathsf{lc}(x)$ $\mathsf{rc}(x)$ context diversity
$\varepsilon$ $\mathtt{{\color[rgb]{1,0,0}\underline{\,}}a{\color[rgb]{1,0,0}\underline{\,}}b{\color[rgb]{1,0,0}\underline{\,}}c{\color[rgb]{1,0,0}\underline{\,}}b{\color[rgb]{1,0,0}\underline{\,}}b{\color[rgb]{1,0,0}\underline{\,}}c{\color[rgb]{1,0,0}\underline{\,}}b{\color[rgb]{1,0,0}\underline{\,}}c{\color[rgb]{1,0,0}\underline{\,}}a{\color[rgb]{1,0,0}\underline{\,}}b{\color[rgb]{1,0,0}\underline{\,}}c{\color[rgb]{1,0,0}\underline{\,}}\mathdollar\,}$ $[1..12]$ $\{\mathtt{\mathdollar},\mathtt{a},\mathtt{b},\mathtt{c}\}$ $\{\mathtt{\mathdollar},\mathtt{a},\mathtt{b},\mathtt{c}\}$ $(4,4)$
$\mathtt{a}$ $\mathtt{\,{\color[rgb]{1,0,0}\underline{a}}\,b\,c\,b\,b\,c\,b\,c\,{\color[rgb]{1,0,0}\underline{a}}\,b\,c\,\mathdollar\,}$ $\{1,9\}$ $\{\mathtt{\mathdollar},\mathtt{c}\}$ $\{\mathtt{b}\}$ $(2,1)$
$\mathtt{b}$ $\mathtt{\,a\,{\color[rgb]{1,0,0}\underline{b}}\,c\,{\color[rgb]{1,0,0}\underline{b}}\,{\color[rgb]{1,0,0}\underline{b}}\,c\,{\color[rgb]{1,0,0}\underline{b}}\,c\,a\,{\color[rgb]{1,0,0}\underline{b}}\,c\,\mathdollar\,}$ $\{2,4,5,7,10\}$ $\{\mathtt{a},\mathtt{b},\mathtt{c}\}$ $\{\mathtt{b},\mathtt{c}\}$ $(3,2)$
$\mathtt{c}$ $\mathtt{\,a\,b\,{\color[rgb]{1,0,0}\underline{c}}\,b\,b\,{\color[rgb]{1,0,0}\underline{c}}\,b\,{\color[rgb]{1,0,0}\underline{c}}\,a\,b\,{\color[rgb]{1,0,0}\underline{c}}\,\mathdollar\,}$ $\{3,6,8,11\}$ $\{\mathtt{b}\}$ $\{\mathtt{\mathdollar},\mathtt{a},\mathtt{b}\}$ $(1,3)$
$\mathtt{ab}$ $\mathtt{\,{\color[rgb]{1,0,0}\underline{a\,b}}\,c\,b\,b\,c\,b\,c\,{\color[rgb]{1,0,0}\underline{a\,b}}\,c\,\mathdollar\,}$ $\{1,9\}$ $\{\mathtt{\mathdollar},\mathtt{c}\}$ $\{\mathtt{c}\}$ $(2,1)$
$\mathtt{bc}$ $\mathtt{\,a\,{\color[rgb]{1,0,0}\underline{b\,c}}\,b\,{\color[rgb]{1,0,0}\underline{b\,c}}\,{\color[rgb]{1,0,0}\underline{b\,c}}\,a\,{\color[rgb]{1,0,0}\underline{b\,c}}\,\mathdollar\,}$ $\{2,5,7,10\}$ $\{\mathtt{a},\mathtt{b},\mathtt{c}\}$ $\{\mathtt{\mathdollar},\mathtt{a},\mathtt{b}\}$ $(3,3)$
$\mathtt{cb}$ $\mathtt{\,a\,b\,{\color[rgb]{1,0,0}\underline{c\,b}}\,b\,{\color[rgb]{1,0,0}\underline{c\,b}}\,c\,a\,b\,c\,\mathdollar\,}$ $\{3,6\}$ $\{\mathtt{b}\}$ $\{\mathtt{b},\mathtt{c}\}$ $(1,2)$
$\mathtt{abc}$ $\mathtt{\,{\color[rgb]{1,0,0}\underline{a\,b\,c}}\,b\,b\,c\,b\,c\,{\color[rgb]{1,0,0}\underline{a\,b\,c}}\,\mathdollar\,}$ $\{1,9\}$ $\{\mathtt{\mathdollar},\mathtt{c}\}$ $\{\mathtt{\mathdollar},\mathtt{b}\}$ $(2,2)$
$\mathtt{bcb}$ $\mathtt{\,a\,{\color[rgb]{1,0,0}\underline{b\,c\,b}}\,{\color[rgb]{1,0,0}\underline{b\,c\,b}}\,c\,a\,b\,c\,\mathdollar\,}$ $\{2,5\}$ $\{\mathtt{a},\mathtt{b}\}$ $\{\mathtt{b},\mathtt{c}\}$ $(2,2)$
$x\in\mathcal{S}^{\mathsf{r}}$ net occurrences $\mathsf{NOcc}(x)$ $\mathsf{NF}(x)$ LMR RMR MR NSMR SMR
$\varepsilon$ $\mathtt{\,a\,b\,c\,b\,b\,c\,b\,c\,a\,b\,c\,\mathdollar\,} \emptyset$ 0 $\surd$ $\surd$ $\surd$
$\mathtt{a}$ $\mathtt{\,a\,b\,c\,b\,b\,c\,b\,c\,a\,b\,c\,\mathdollar\,} \emptyset$ 0 $\surd$
$\mathtt{b}$ $\mathtt{\,a\,b\,c\,b\,b\,c\,b\,c\,a\,b\,c\,\mathdollar\,} \emptyset$ 0 $\surd$ $\surd$ $\surd$
$\mathtt{c}$ $\mathtt{\,a\,b\,c\,b\,b\,c\,b\,c\,a\,b\,c\,\mathdollar\,} \emptyset$ 0 $\surd$
$\mathtt{ab}$ $\mathtt{\,a\,b\,c\,b\,b\,c\,b\,c\,a\,b\,c\,\mathdollar\,} \emptyset$ 0 $\surd$
$\mathtt{bc}$ $\mathtt{\,a\,b\,c\,b\,b\,c\,{\color[rgb]{1,0,0}\underline{b\,c}}\,a\,b\,c\,\mathdollar\,}$ $\{7\}$ 1 $\surd$ $\surd$ $\surd$ $\surd$
$\mathtt{cb}$ $\mathtt{\,a\,b\,c\,b\,b\,c\,b\,c\,a\,b\,c\,\mathdollar\,} \emptyset$ 0 $\surd$
$\mathtt{abc}$ $\mathtt{\,{\color[rgb]{1,0,0}\underline{a\,b\,c}}\,b\,b\,c\,b\,c\,{\color[rgb]{1,0,0}\underline{a\,b\,c}}\,\mathdollar\,}$ $\{1,9\}$ 2 $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
$\mathtt{bcb}$ $\mathtt{\,a\,{\color[rgb]{1,0,0}\underline{b\,c\,b}}\,{\color[rgb]{1,0,0}\underline{b\,c\,b}}\,c\,a\,b\,c\,\mathdollar\,}$ $\{2,5\}$ 2 $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
Table 1: An example of the concepts introduced in Section˜2.3 for $T=\mathtt{abcbbcbcabc\mathdollar}$. For every repeat $x\in\mathcal{S}^{\mathsf{r}}$, we show a visualization of the occurrences of $x$, $\mathsf{Occ}(x)$, $\mathsf{lc}(x)$, $\mathsf{rc}(x)$, and the context diversity of $x$ in the upper table. In the lower table, we show a visualization of the net occurrences of $x$, $\mathsf{NOcc}(x)$, $\mathsf{NF}(x)$, and subsequently check if $x$ is in each of $\mathcal{S}^{\mathsf{lmr}}$, $\mathcal{S}^{\mathsf{rmr}}$, $\mathcal{S}^{\mathsf{mr}}$, $\mathcal{S}^{\mathsf{nsmr}}$ and $\mathcal{S}^{\mathsf{smr}}$.
2.4 Suffix trees and suffix arrays
The suffix trie of $T$ is the tree defined on the set $\mathcal{S}$ of nodes and edges from $x\in\mathcal{S}$ to $xc\in\mathcal{S}$ with $c\in\Sigma_{\mathdollar}$. The suffix tree of $T$ is the compacted suffix trie in which every non-branching internal node of the suffix trie is removed and considered to be an implicit node. Every internal node of the suffix tree is a right-maximal repeat $x$ and has $|\mathsf{rc}(x)|$ children. A link from a node $x$ to a (possibly implicit) node $ax\in\mathcal{S}$ for some $a\in\Sigma_{\mathdollar}$ is called a Weiner link. It is known that every node can be reached by following Weiner links from the root $\varepsilon$, and the total number of Weiner links is at most $3n$, which can be seen from the duality between suffix trees and DAWGs [5].
The suffix array $\mathsf{SA}[1..n]$ of $T$ is the integer array such that, for any $i\in[1..n]$, $T[\mathsf{SA}[i]..]$ is the lexicographically $i$-th suffix among the non-empty suffixes of $T$. For any string $x\in\mathcal{S}$, the SA-interval of $x$, denoted by $\mathcal{I}(x)$, is the maximal interval $[p..q]$ such that $x$ is a prefix of $T[\mathsf{SA}[i]..]$ for any $i\in[p..q]$. Then, it holds that $\mathsf{Occ}(x)=\{\mathsf{SA}[i]\mid i\in\mathcal{I}(x)\}$.
2.5 Burrows-Wheeler transform
The Burrows-Wheeler transform (BWT) $\mathsf{L}[1..n]$ of $T$ is the string such that $\mathsf{L}[i]=T[\mathsf{SA}[i]-1]$ for any $i\in[1..n]$. By definition, it holds that $\mathsf{lc}(x)=\{\mathsf{L}[i]\mid i\in\mathcal{I}(x)\}$ for any string $x$. The LF-mapping $\mathsf{LF}$ is the permutation on $[1..n]$ such that $\mathsf{LF}(i)$ represents the lexicographic rank of the suffix $T[\mathsf{SA}[i]-1..]$ if $\mathsf{SA}[i]\neq 1$, and otherwise $\mathsf{LF}(i)=1$. The FL-mapping $\mathsf{FL}$ is the inverse mapping of $\mathsf{LF}$. Given an integer $i$ in $\mathcal{I}(x)$ for a string $x$, we can compute $x$ by using FL-mapping $|x|$ times from $i$ because $x[k]=\mathsf{L}[\mathsf{FL}^{k}(i)]$ holds for any $k~(1\leq k\leq|x|)$, where $\mathsf{FL}^{k}(i)$ is defined recursively $\mathsf{FL}^{k}(i)=\mathsf{FL}(\mathsf{FL}^{k-1}(i))$ with the base case $\mathsf{FL}^{1}(i)=\mathsf{FL}(i)$.
Table˜2 shows an example of $\mathsf{SA}$, $\mathsf{LF}(\cdot)$, $\mathsf{FL}(\cdot)$ and $\mathsf{L}$ for $T=\mathtt{abcbbcbcabc\mathdollar}$. Figure˜1 also illustrates the suffix tree of $T=\mathtt{abcbbcbcabc\mathdollar}$ and a relationship between Weiner links and BWTs.
$i$ $\mathsf{SA}[i]$ $\mathsf{LF}(i)$ $\mathsf{FL}(i)$ $\mathsf{L}[i]$ $T[\mathsf{SA}[i]..]$
[…]
2 9 10 5 $\mathtt{c}$ $\mathtt{abc\mathdollar}$
[…]
4 4 11 8 $\mathtt{c}$ $\mathtt{bbcbcabc\mathdollar}$
5 10 2 9 $\mathtt{a}$ $\mathtt{bc\mathdollar}$
6 7 12 10 $\mathtt{c}$ $\mathtt{bcabc\mathdollar}$
7 2 3 11 $\mathtt{a}$ $\mathtt{bcbbcbcabc\mathdollar}$
8 5 4 12 $\mathtt{b}$ $\mathtt{bcbcabc\mathdollar}$
9 11 5 1 $\mathtt{b}$ $\mathtt{c\mathdollar}$
10 8 6 2 $\mathtt{b}$ $\mathtt{cabc\mathdollar}$
11 3 7 4 $\mathtt{b}$ $\mathtt{cbbcbcabc\mathdollar}$
12 6 8 6 $\mathtt{b}$ $\mathtt{cbcabc\mathdollar}$
Table 2: An example of $\mathsf{SA}[i]$, $\mathsf{LF}(i)$, $\mathsf{FL}(i)$ and $\mathsf{L}[i]$ for $T=\mathtt{abcbbcbcabc\mathdollar}$.
Figure 1: The left figure shows the suffix tree of $T=\mathtt{abcbbcbcabc\mathdollar}$ illustrated over sorted suffixes, on which each node $x$ can be represented by $\mathcal{I}(x)$ and $|x|$. A solid box is a node (highlighted for internal nodes), and a dotted box is an implicit node. The right figure shows the Weiner links outgoing from all internal nodes (the Weiner links from leaves are omitted). The Weiner links that points to internal nodes are depicted with solid arrows, and the other ones with dotted arrows. Observe that there is a Weiner link from a node $x$ to axax for any character $a\in\mathsf{lc}(x)=\{\mathsf{L}[i]\mid i\in\mathcal{I}(x)\}$, where the case with $a=\mathdollar$ is excluded unless $x=\varepsilon$.
The run-length encoded BWT (RLBWT) of $T$ represents $\mathsf{L}$ in $O(r)$ space by string $\mathsf{L}^{\prime}[1..r]\in\Sigma_{\mathdollar}^{r}$ and array $\mathsf{d}[1..r]\in[1..n]^{r}$ such that $\mathsf{L}=\mathsf{L}^{\prime}[1]^{\mathsf{d}[1]}\mathsf{L}^{\prime}[2]^{\mathsf{d}[2]}\dots\mathsf{L}^{\prime}[r]^{\mathsf{d}[r]}$ and $\mathsf{L}^{\prime}[i]\neq\mathsf{L}^{\prime}[j]$ for any $1\leq i<j\leq r$. Each component $\mathsf{L}^{\prime}[i]^{\mathsf{d}[i]}$ is called a run of $\mathsf{L}$. While $r\leq n$ is obvious, inequality $\sigma\leq r$ holds under the assumption that $T$ contains all characters in $\Sigma_{\mathdollar}$.
Example 4.
For $T=\mathtt{abcbbcbcabc\mathdollar}$, its BWT $\mathsf{L}=\mathtt{cc\mathdollar cacabbbbb}$ has $r=7$ runs. Then, we have $\mathsf{L}^{\prime}=\mathtt{c\mathdollar cacab}$ and $\mathsf{d}=[2,1,1,1,1,1,5]$.
A notable property of $\mathsf{LF}$ is that $\mathsf{LF}(i)=k_{1}+k_{2}$ holds, where $k_{1}$ is the number of characters smaller than $\mathsf{L}[i]$ in $\mathsf{L}$ and $k_{2}$ is the number of occurrences of $\mathsf{L}[i]$ in $\mathsf{L}[1..i]$. In particular, if $\mathsf{L}[p..q]$ consists of a single character, then $\mathsf{LF}(i)=\mathsf{LF}(p)+i-p$ for any $i\in[p..q]$, and $\mathsf{FL}(i^{\prime})=\mathsf{FL}(p^{\prime})+i^{\prime}-p^{\prime}$ for any $i^{\prime}\in[p^{\prime}..q^{\prime}]$ with $p^{\prime}=\mathsf{LF}(p)$ and $q^{\prime}=\mathsf{LF}(q)$. Hence, LF-mapping can be implemented in $O(r)$ space by constructing the predecessor data structure for the set of beginning positions of runs of $\mathsf{L}$ and remembering the $\mathsf{LF}(\cdot)$ values for those positions (FL-mapping can be implemented similarly). Using the predecessor data structure of Lemma˜2, we get:
Lemma 5.
Given the RLBWT of $T$, we can construct an $O(r)$-size data structure in $O(r\log\log_{w}(n/r))$ time and $O(r)$ space to compute $\mathsf{LF}(i)$ and $\mathsf{FL}(i)$ in $O(\log\log_{w}(n/r))$ time for any $i\in[1..n]$.
Nishimoto and Tabei [24] proposed a novel data structure, called the move data structure, to implement LF-mapping by a simple linear search on a list of intervals without predecessor queries. We consider a list $p_{1}..p_{2}),[p_{2}..p_{3}),\dots,[p_{\hat{r}}..p_{\hat{r}+1})$ of $\hat{r}=O(r)$ intervals such that $p_{1}=1$, $p_{\hat{r}+1}=n+1$, and for any $k\in[1..\hat{r}$], $\mathsf{L}[p_{k}..p_{k+1})$ consists of a single character and $\mathsf{LF}(p_{k})..\mathsf{LF}(p_{k+1}-1)]$ intersects with $O(1)$ intervals in the list. We let each $p_{k}..p_{k+1})$ have a pointer to the interval $p_{k^{\prime}}..p_{k^{\prime}+1})$ that contains $\mathsf{LF}(p_{k})$ and offset $o_{k}=\mathsf{LF}(p_{k})-p_{k^{\prime}}$. Then, $\mathsf{LF}(i)=\mathsf{LF}(p_{k})+i-p_{k}=p_{k^{\prime}}+o_{k}+i-p_{k}$ holds for any $i\in[p_{k}..p_{k+1})$. Given $i$ and interval $p_{k}..p_{k+1})$ that contains $i$, we can compute $\mathsf{LF}(i)$ and find the interval that contains $\mathsf{LF}(i)$ by assessing $O(1)$ intervals from $p_{k^{\prime}}..p_{k^{\prime}+1})$ to the right in the list. Note that $\mathsf{LF}(i)$ computation on the move data structure works along with the pointer to the interval in the list that contains $i$, which is handled implicitly hereafter. Under the assumption that $\sigma\leq r$ in this paper, we can construct the data structure needed for LF-mapping (and FL-mapping) in $O(r\log r)$ time [25, Appendix D]:
Lemma 6.
Given the RLBWT of $T$, we can construct an $O(r)$-size data structure in $O(r\log r)$ time and $O(r)$ space to compute $\mathsf{LF}(i)$ and $\mathsf{FL}(i)$ in $O(1)$ time for any $i\in[1..n]$.
3 Upper bound $2r$ of net occurrences and MUSs
In this section, we consider combinatorial properties of net occurrences on BWTs.
We start with the following lemma.
Lemma 7.
For any integer $i\in[p..q]=\mathcal{I}(x)$ for a repeat $x$ of $T$, $\mathsf{SA}[i]$ is a net occurrence of $x$ if and only if $\mathsf{L}[i]$ is unique in $\mathsf{L}[p..q]$ and $T[\mathsf{SA}[i]..\mathsf{SA}[i]+|x|]$ is unique in $T$.
Proof.
If $a=\mathsf{L}[i]$ is not unique in $\mathsf{L}[p..q]$, then $\mathsf{SA}[i]$ is not a net occurrence of $x$ because $ax=T[\mathsf{SA}[i]-1..\mathsf{SA}[i]+|x|)$ is a repeat that covers $\mathsf{SA}[i]..\mathsf{SA}[i]+|x|)$. Similarly, if $T[\mathsf{SA}[i]..\mathsf{SA}[i]+|x|]$ is not unique in $T$, then $\mathsf{SA}[i]$ is not a net occurrence of $x$ because $T[\mathsf{SA}[i]..\mathsf{SA}[i]+|x|]$ is a repeat that covers $\mathsf{SA}[i]..\mathsf{SA}[i]+|x|)$.
On the other hand, if $\mathsf{L}[i]$ is unique in $\mathsf{L}[p..q]$ and $T[\mathsf{SA}[i]..\mathsf{SA}[i]+|x|]$ is unique in $T$, there is no repeat that covers $\mathsf{SA}[i]..\mathsf{SA}[i]+|x|)$, and hence, $\mathsf{SA}[i]$ is a net occurrence of $x$. ∎
Example 8.
Let us consider a repeat $\mathtt{bc}$ with $\mathcal{I}(\mathtt{bc})=[5..8]$ in the example of Table˜2. $\mathsf{SA}[6]=7$ is a net occurrence of $\mathtt{bc}$ because $\mathsf{L}[6]=\mathtt{c}$ is unique in $\mathsf{L}[5..8]=\mathtt{acab}$, and $T[\mathsf{SA}[6]..\mathsf{SA}[6]+2]=T[7..9]=\mathtt{bca}$ is unique in $T$. $\mathsf{SA}[5]=10$ and $\mathsf{SA}[7]=2$ are not net occurrences of $\mathtt{bc}$ because $\mathsf{L}[5]=\mathsf{L}[7]=\mathtt{a}$ is not a unique character in $\mathsf{L}[5..8]$. $\mathsf{SA}[7]=2$ and $\mathsf{SA}[8]=5$ are not net occurrences of $\mathtt{bc}$ because $T[2..4]=T[5..7]=\mathtt{bcb}$ is a repeat.
The following lemma was implicitly proved in [10, Theorem 24] to show the total length of all net occurrences is bounded by twice the sum of irreducible LCPs.
Lemma 9.
If $\mathsf{SA}[i]$ is a net occurrence of a repeat $x$ of $T$, then $i$ is at the beginning or ending position of a run in $\mathsf{L}$, and $\mathsf{SA}[i]$ cannot be a net occurrence of another repeat.
Proof.
Let $p..q]=\mathcal{I}(x)$. If $i\in[p..q]$ is not at the run’s boundaries of $\mathsf{L}$, then at least one of $\mathsf{L}[i]=\mathsf{L}[i+1]$ with $i+1\in[p..q]$ or $\mathsf{L}[i-1]=\mathsf{L}[i]$ with $i-1\in[p..q]$ holds, which means that $\mathsf{SA}[i]$ is not a net occurrence of $x$ by Lemma˜7. $\mathsf{SA}[i]$ cannot be a net occurrence of two distinct repeats, since otherwise, the occurrence of the shorter repeat is covered by the longer one. ∎
Lemma˜9 leads to the following theorem:
Theorem 10.
The total number of net occurrences of $T$ is less than $2r$.
Proof.
Let $r_{1}$ and respectively $r_{>1}$ be the number of runs in $\mathsf{L}$ of length $1$ and longer than $1$, i.e., $r=r_{1}+r_{>1}$. Since there are $r_{1}+2r_{>1}$ positions that correspond to beginning or ending positions of runs in $\mathsf{L}$, it is clear from Lemma˜9 that the total number of net occurrences is at most $r_{1}+2r_{>1}\leq 2r$. The number is shown to be less than $2r$ by looking closer at some special positions like $1$ and $n$ on $\mathsf{L}$: Since $\mathsf{SA}[1]$ (resp. $\mathsf{SA}[n]$) cannot be a net occurrence if the first (resp. last) run of $\mathsf{L}$ is longer than $1$, only one net occurrence is charged to the first (resp. last) run of $\mathsf{L}$. ∎
A non-empty interval $[b..e]$ is called a minimal unique substring (MUS) if $T[b..e]$ is unique while $T[b+1..e]$ and T[b..e1]T[b..e-1] are repeats. In [14, 21], it has been proved that the net occurrences and MUSs are dual concepts, basically saying that for any two consecutive MUSs $[b..e]$ and $b^{\prime}..e^{\prime}]$ with $b<b^{\prime}$, $b+1$ is a net occurrence of $T[b+1..e^{\prime}-1]$. With this duality, we immediately get the following:
Corollary 11.
The number of MUSs of $T$ is less than $2r$.
4 R-enum revisited
4.1 Rough sketch of r-enum and speedup
Conceptually, r-enum performs a breadth-first traversal on the internal nodes $\mathcal{S}^{\mathsf{rmr}}$ of the suffix tree by following Weiner links using RLBWT-based data structures. Each node $x\in\mathcal{S}^{\mathsf{rmr}}$ is visited with its suffix interval $\mathcal{I}(x)$ and length $|x|$ together with the information about its right extensions $\mathsf{rlist}(x)$, which is the list of the set $\{(c,\mathcal{I}(xc))\mid c\in\mathsf{rc}(x)\}$ sorted by the character $c$. The triplet $\mathsf{repr}(x)=(\mathcal{I}(x),\mathsf{rlist}(x),|x|)$ is called the rich representation [2, 26] of $x$, which can be stored in $O(|\mathsf{rc}(x)|)$ space.
Given $\mathsf{repr}(x)=([p..q],\{(c_{i},[p_{i}..q_{i}])\}_{i=1}^{k},|x|)$, we can compute $\{\mathsf{repr}(ax)\mid ax\in\mathcal{S},a\in\Sigma\}$ as follows: A null-initialized array $A[1..\sigma]$ is used as a working space to build $\mathsf{rlist}(ax)$ at $A[a]$ for $a\in\mathsf{lc}(x)\setminus\{\mathdollar\}$. For every $i\in[1..k]$, we compute $\mathsf{lc}(xc_{i})$ by a range distinct query on $\mathsf{L}[p_{i}..q_{i}]$, and for each $a\in\mathsf{lc}(xc_{i})\setminus\{\mathdollar\}$ we compute $\mathcal{I}(axc_{i})$ by LF-mapping and append $(c_{i},\mathcal{I}(axc_{i}))$ to the list being built at $A[a]$. After processing all $i\in[1..k]$, $A[a]$ becomes $\mathsf{rlist}(ax)$. Finally, we enumerate $\mathsf{lc}(x)$ by a range distinct query on $\mathsf{L}[p..q]$ to go through only the used entries of $A$ to compute $\mathsf{repr}(ax)$ and clear the entries. For the next round of the breadth-first traversal on $\mathcal{S}^{\mathsf{rmr}}$, we keep only the set $\{\mathsf{repr}(ax)\mid ax\in\mathcal{S}^{\mathsf{rmr}}\}$, discarding $\mathsf{repr}(ax)$ if $\mathsf{rlist}(ax)$ contains only one element.
Example 12.
In the r-enum for $T=\mathtt{abcbbcbcabc\mathdollar}$, we visit a repeat $\mathtt{bc}\in\mathcal{S}^{\mathsf{rmr}}$ with its rich representation $\mathsf{repr}(\mathtt{bc})=(\mathcal{I}(\mathtt{bc}),\mathsf{rlist}(\mathtt{bc}),|\mathtt{bc}|)=([5..8],\{(\mathdollar,[5..5]),(\mathtt{a},[6..6]),(\mathtt{b},[7..8])\},2)$. With a null-initialized array $A[1..\sigma]$, we compute $\mathsf{repr}(\mathtt{abc})=([2..3],\{(\mathdollar,[2..2]),(\mathtt{b},[3..3])\},3)$ at $A[\mathtt{a}]$, $\mathsf{repr}(\mathtt{bbc})=([4..4],\{(\mathtt{b},[4..4])\},3)$ at $A[\mathtt{b}]$, and $\mathsf{repr}(\mathtt{cbc})=([12..12],\{(\mathtt{a},[12..12])\},3)$ at $A[\mathtt{c}]$.
R-enum executes the above procedure in $O(r)$ space using range distinct queries on $\mathsf{L}^{\prime}[1..r]$ and RLBWT-based LF-mapping. The breadth-first traversal on $\mathcal{S}^{\mathsf{rmr}}$ can be made in $O(r)$ space because the space needed to store the rich representations for all strings in $\{ax\in\mathcal{S}\mid a\in\Sigma,x\in\mathcal{S}^{\mathsf{rmr}},|x|=t\}$ for any $t\in[0..n]$ was proved to be $O(r)$ in Section 3.3 of [26]. The total number of outputs of range distinct queries is upper bounded by $O(n)$ because each output can be charged to a distinct Weiner link. Similarly, the total number of LF-mapping needed is bounded by $O(n)$.
The original r-enum used Lemma˜5 to implement LF-mapping and runs in $O(n\log\log_{w}(n/r))$ time in total (excluding the time to output patterns for MAWs). We show that this can be improved to $O(n)$ time.
Theorem 13.
Given the RLBWT of $T$, r-enum runs in $O(r)$ working space and $O(n)$ time in addition to the time linear to the number of output strings.
Proof.
If $r=\Omega(n/\log n)$, the original r-enum with Lemma˜5 is already upper bounded by $O(n)$ because $O(n\log\log_{w}(n/r))=O(n\log\log_{w}\log n)$ and $w=\Theta(\log n)$. For the case with $r=o(n/\log n)$, we use the move data structure of Lemma˜6 to implement LF-mapping, which can be constructed in $O(r\log r)=O(n)$ time. Since LF-mapping can be done in $O(1)$ time on the move data structure, the r-enum runs in $O(n)$ time. ∎
4.2 Extension for NSMRs, SMRs, and context diversities
In this subsection, we extend r-enum to compute the NSMRs and SMRs, and the context diversity for every RMR.
Recall that r-enum traverses $x\in\mathcal{S}^{\mathsf{rmr}}$ while computing $\{\mathsf{repr}(ax)\mid ax\in\mathcal{S}\}$ from $\mathsf{repr}(x)=([p..q],\{(c_{i},[p_{i}..q_{i}])\}_{i=1}^{k},|x|)$. First, we can compute the context diversity $(|\mathsf{lc}(x)|,|\mathsf{rc}(x)|)$ of $x\in\mathcal{S}^{\mathsf{rmr}}$ because $|\mathsf{rc}(x)|$ is $k$ and $|\mathsf{lc}(x)|$ is the number of outputs of a range distinct query on $\mathsf{L}[p..q]$. Since $x$ is an SMR if and only if $|\mathsf{lc}(x)|=|\mathsf{rc}(x)|=|\mathsf{Occ}(x)|=q-p+1$, the context diversity has enough information to choose SMRs from RMRs. On the other hand, NSMRs cannot be determined from the context diversities.
[…]
Since we can check the conditions for NSMRs and SMRs along with the traversal of $\mathcal{S}^{\mathsf{rmr}}$, the time and space complexities of r-enum are retained.
Theorem 14.
Given the RLBWT of $T$, we can compute the NSMRs and SMRs, and the context diversity for every RMR in $O(n)$ time and $O(r)$ working space. The net occurrences of output patterns can also be computed in the same time and space complexities.
5 An $O(r)$-size data structure for NF-queries
In this section, we tackle the problem called Single-NF [10] on RLBWTs: We consider preprocessing the RLBWT of $T$ to construct a data structure of size $O(r)$ so that we can support NF-queries of computing $\mathsf{NF}(P)$ for any query pattern $P$.
We define the reversed trie for a set $S$ of strings as follows:
  • The nodes consist of the suffixes of strings in $S$.
  • There is an edge from node $x$ to $ax$ with a character $a$.
Theorem 15.
Given the RLBWT of $T$, we can build an $O(r)$-size data structure in $O(r)$ space and in $O(n)$ time with high probability or deterministic $O(n+|\mathcal{S}^{\mathsf{nsmr}}|\log\log\min(\sigma,|\mathcal{S}^{\mathsf{nsmr}}|))$ time to support NF-queries for any query pattern $P$ in $O(|P|)$ time.
Proof.
We execute r-enum of Theorem˜14, which actually performs a breadth-first traversal of the reversed trie for $\mathcal{S}^{\mathsf{rmr}}$ in $O(n)$ time and $O(r)$ space. Our idea is to build the compacted reversed trie $\mathcal{T}$ for $\mathcal{S}^{\mathsf{nsmr}}$ during the traversal, which can be stored in $O(|\mathcal{S}^{\mathsf{nsmr}}|)=O(r)$ space because $|\mathcal{S}^{\mathsf{nsmr}}|\leq 2r$ by Lemma˜9. For each node $x$ of $\mathcal{T}$, which corresponds to a string in $\mathcal{S}^{\mathsf{nsmr}}$ or a branching node, we store the string length $|x|$ and an integer $i_{x}$ in $\mathcal{I}(x)$ so that the edge label between $x$ and its parent $y$ can be retrieved by using FL-mapping |x||y||x|-|y| times from $i_{x}$. For a node $x\in\mathcal{S}^{\mathsf{nsmr}}$, we also store the net frequency of $x$. See Figure˜2 for an illustration of $\mathcal{T}$.
We show that we can build $\mathcal{T}$ in $O(n)$ time and $O(r)$ working space during the breadth-first traversal of r-enum. For each string depth $t$, we maintain the compacted reversed trie $\mathcal{T}_{t}$ for the set $\mathcal{S}^{\mathsf{rmr}}_{t}\cup\mathcal{S}^{\mathsf{nsmr}}_{\leq t}$, where $\mathcal{S}^{\mathsf{rmr}}_{t}$ is the set of RMRs of length $t$ and $\mathcal{S}^{\mathsf{nsmr}}_{\leq t}$ is the set of NSMRs of length $\leq t$. Since $|\mathcal{S}^{\mathsf{rmr}}_{t}|=O(r)$ and $|\mathcal{S}^{\mathsf{nsmr}}_{\leq t}|=O(r)$, the size of $\mathcal{T}_{t}$ is $O(r)$. In the next round of the breadth-first traversal, r-enum extends every string in $\mathcal{S}^{\mathsf{rmr}}_{t}$ by one character to the left to get $\mathcal{S}^{\mathsf{rmr}}_{t+1}$. We update $\mathcal{T}_{t}$ to $\mathcal{T}_{t+1}$ by processing $x\in\mathcal{S}^{\mathsf{rmr}}_{t}$ as follows, where $A_{x}=\{ax\mid ax\in\mathcal{S}^{\mathsf{rmr}}_{t+1}\}$:
  • If $|A_{x}|\geq 2$, we add new edges from $x$ to the strings in $A_{x}$.
  • If $|A_{x}|=1$, we add a new edge from $x$ to the string in $A_{x}$, and remove the non-branching node $x$ if $x$ is not an NSMR.
  • If $|A_{x}|=0$ and $x\notin\mathcal{S}^{\mathsf{nsmr}}$, we remove the node $x$ and the edge to its parent $y$. We also remove $y$, if it becomes a non-branching node not in $\mathcal{S}^{\mathsf{nsmr}}$.
In total, these tasks can be done in $O(n)$ time and $O(r)$ space along with r-enum of Theorem˜14. For each branching node of $\mathcal{T}$ with $k\leq\min(\sigma,|\mathcal{S}^{\mathsf{nsmr}}|)$ outgoing edges, we build a dictionary of Lemma˜1 in $O(k)$ space and $O(k)$ time w.h.p. or deterministic $O(k\log\log\min(\sigma,k))$ time so that we can decide which child to proceed by the first characters of edges in $O(1)$ time. The construction time of the dictionaries for all branching nodes is bounded by $O(|\mathcal{S}^{\mathsf{nsmr}}|)$ time w.h.p. or deterministic $O(|\mathcal{S}^{\mathsf{nsmr}}|\log\log\min(\sigma,|\mathcal{S}^{\mathsf{nsmr}}|))$ time.
Given a query pattern $P$, we read the characters of $P$ from right to left and traverse $\mathcal{T}$ from the root in $O(|P|)$ time. If we can reach a node $P\in\mathcal{S}^{\mathsf{nsmr}}$, we output $\mathsf{NF}(P)$ stored in the node. ∎
We can also support queries of computing the net occurrences of a query pattern $P$ in $O(|P|+|\mathsf{NOcc}(P)|)$ time by storing $\mathsf{NOcc}(x)$ for each node $x\in\mathcal{S}^{\mathsf{nsmr}}$ in $\mathcal{T}$, which can be stored in $O(r)$ space in total due to Lemma˜9.
Figure 2: An illustration of the compacted reversed trie $\mathcal{T}$ for $\mathcal{S}^{\mathsf{nsmr}}=\{\mathtt{bc},\mathtt{abc},\mathtt{bcb}\}$ in our running example $T=\mathtt{abcbbcbcabc\mathdollar}$. A node corresponding to an NSMR is highlighted. Storing $\mathsf{NOcc}(\cdot)$ is optional. Note that edge labels are not stored explicitly. For example, the string $\mathtt{bcb}$ on the edge from $\varepsilon$ to $\mathtt{bcb}$ is retrieved using FL-mapping $i_{\mathtt{bcb}}-i_{\varepsilon}=3$ times from $i_{\mathtt{bcb}}=7$ when necessary.
6 Conclusions
We improved the time complexity of r-enum, and extended its enumeration power for other context-sensitive repeats in the literature. We also proposed the first data structure of size $O(r)$ to support NF-queries in optimal time. As a combinatorial property, we formally proved that the total number of net occurrences, as well as MUSs, is upper bounded by $O(r)$. We finally remark that our work leads to the first algorithm to compute the sorted list of net occurrences and MUSs in $O(n)$ time and $O(r)$ space: We enumerate all net occurrences by Theorem˜13, and sort them in $O(n)$ time and $O(r)$ space by [25, Lemma 13].
Acknowledgements
Tomohiro I was supported by JSPS KAKENHI (Grant Numbers JP24K02899) and JST AIP Acceleration Research JPMJCR24U4, Japan.

### v2

@@
 Given the run-length encoded BWT (RLBWT) of $T$, r-enum runs in $O(n\log\log_{w}(n/r))$ time in addition to the time linear to the number of output strings, where $w=\Theta(\log n)$ is the word size.
-In this paper, we improve the $O(n\log\log_{w}(n/r))$ term to $O(n)$.
-We also extend r-enum to compute other context-sensitive repeats such as near-supermaximal repeats (NSMRs) and supermaximal repeats, as well as the context-diversity for every maximal repeat in the same complexities.
-Furthermore, we study the occurrences that witness NSMRs, which have recently attracted attention under the name of net occurrences: An occurrence of a repeat is called a net occurrence if it is not covered by another repeat, and the net frequency of a repeat is the number of its net occurrences.
-With this terminology, an NSMR is defined to be a repeat with a positive net frequency.
+In this paper, we first improve the $O(n\log\log_{w}(n/r))$ term to $O(n)$.
+We next extend r-enum to compute other context-sensitive repeats such as near-supermaximal repeats (NSMRs) and supermaximal repeats, as well as the context diversity for every maximal repeat in the same complexities.
+Furthermore, we study net occurrences: An occurrence of a repeat is called a net occurrence if it is not covered by another repeat, and the net frequency of a repeat is the number of its net occurrences.
+With this terminology, an NSMR is a repeat with a positive net frequency.
 Given the RLBWT of $T$, we show how to compute the set $\mathcal{S}^{\mathsf{nsmr}}$ of all NSMRs in $T$ together with their net frequency/occurrences in $O(n)$ time and $O(r)$ space.
-We also show that an $O(r)$-space data structure can be built from the RLBWT to support queries of computing the net frequency of any query pattern $P$ in optimal $O(|P|)$ time.
+We also show that an $O(r)$-space data structure can be built from the RLBWT to compute the net frequency/occurrences of any pattern in optimal time.
 The data structure is built in $O(r)$ space and in $O(n)$ time with high probability or deterministic $O(n+|\mathcal{S}^{\mathsf{nsmr}}|\log\log\min(\sigma,|\mathcal{S}^{\mathsf{nsmr}}|))$ time, where $\sigma\leq r$ is the alphabet size of $T$.
@@
 With the duality between net occurrences and minimal unique substrings (MUSs), we get a new upper bound $2r$ of the number of MUSs in $T$, which may be of independent interest.
+keywords:
+Supermaximal repeats; Largest maximal repeats; Net frequencies; Run-length Burrows-Wheeler transform; Compressed data mining
 1 Introduction
@@
 One of the simplest patterns based on frequencies would be a repeat, which occurs at least twice in $T$.
-Since the number of distinct repeats in $T$ of length $n$ can be $\Theta(n^{2})$, maximality is often utilized to reduce the number down to $O(n)$ [13].
-A substring of $T$ is called right-maximal (resp.
-left-maximal) if it cannot be extended to the right (resp.
+Since the number of distinct repeats in $T$ of length $n$ can be $\Theta(n^{2})$, maximality is often utilized to reduce the number down to $O(n)$ [1997Gusfield_AlgorOnStrinTreesAnd].
+A substring of $T$ is called a right-maximal repeat (RMR) (resp.
+left-maximal repeat (LMR)) if it cannot be extended to the right (resp.
 to the left) without reducing its number of occurrences in $T$.
 If a repeat is both right-maximal and left-maximal, it is called a maximal repeat (MR).
-Gusfield [13] also introduced stricter criteria to reduce the redundancy in maximal repeats.
+Gusfield [1997Gusfield_AlgorOnStrinTreesAnd] also introduced stricter criteria to reduce the redundancy in maximal repeats.
 A repeat is called a supermaximal repeat (SMR) if it is not a substring of another repeat, and a near-supermaximal repeat (NSMR) if at least one of its occurrences is not covered by another repeat.
 It is easy to see that an SMR is an NSMR, and an NSMR is an MR.
-NSMRs were studied under the name of Chinese frequent strings in [17, 18, 28], and largest maximal repeats in [23, 7].
-Gallé and Tealdi [8, 9] proposed the concept of context diversities to define a general class of context-sensitive repeats including MRs and SMRs.
-Enumerating these context-sensitive repeats has numerous applications in bioinformatics [13, 23, 1] and natural language processing [17, 18, 29, 20, 7] especially valuable for languages with no clear word segmentation such as Chinese and Japanese.
-Motivated by these applications, efficient enumeration algorithms have been extensively studied [29, 1, 27, 2, 26] using a family of suffix data structures such as suffix trees [31], directed acyclic word graphs (DAWGs) [5], compact DAWGs (CDAWGs) [4, 16], suffix arrays [19] and Burrows-Wheeler transform (BWT) [6].
-In this paper, we focus on r-enum [26], an algorithm to enumerate the MRs, minimal unique substrings (MUSs) [14] and minimal absent words (MAWs) in $O(r)$ words of compressed working space, where $r\leq n$ is the number of runs in the BWT of $T$ of length $n$.
+NSMRs were studied under the name of Chinese frequent strings in [2001LinY_ExtracChinesFrequenStrinWithout, 2004LinY_ProperAndFurthApplicOf, 2024OhlebuschBO_FasterComputOfChinesFrequen_SPIRE], and largest maximal repeats in [2008NicolasRSPCDTVM_ModelLocalRepeatOnGenom, 2013Galle_BagOfRepeatRepresOf_SIGIR].
+Gallé and Tealdi [2014GalleT_ContexDiverRepeatAndTheir_LATA, 2018GalleT_EmphxRepeatNewTaxonOf] proposed the concept of context diversities to define a general class of context-sensitive repeats including MRs and SMRs.
+Enumerating these context-sensitive repeats has numerous applications in bioinformatics [1997Gusfield_AlgorOnStrinTreesAnd, 2008NicolasRSPCDTVM_ModelLocalRepeatOnGenom, 2009BecherDH_EfficComputOfAllPerfec] and natural language processing [2001LinY_ExtracChinesFrequenStrinWithout, 2004LinY_ProperAndFurthApplicOf, 2009OkanoharaT_TextCategWithAllSubst_SDM, 2011MasadaTSO_ClustDocumWithMaximSubst, 2013Galle_BagOfRepeatRepresOf_SIGIR] especially valuable for languages with no clear word segmentation such as Chinese and Japanese.
+Motivated by these applications, efficient enumeration algorithms have been extensively studied [2009OkanoharaT_TextCategWithAllSubst_SDM, 2009BecherDH_EfficComputOfAllPerfec, 2015OhlebuschB_AlphabIndepAlgorForFindin, 2020BelazzouguiCKM_LinearTimeStrinIndexAnd, 2021NishimotoT_REnumEnumerOfCharac_CPM] using a family of suffix data structures such as suffix trees [1973Weiner_LinearPatterMatchAlgor], directed acyclic word graphs (DAWGs) [1985BlumerBHECS_SmallAutomRecogSubworOf], compact DAWGs (CDAWGs) [1987BlumerBHME_ComplInverFilesForEffic, 2005InenagaHSTAMP_LineConstOfCompacDirec], suffix arrays [1993ManberM_SuffixArrayNewMethodFor, JDA2004AbouelhodaKO_ReplacSuffixTreesWithEnhan] and Burrows-Wheeler transform (BWT) [1994BurrowsW_BlockSortinLosslDataCompr].
+In this paper, we focus on r-enum [2021NishimotoT_REnumEnumerOfCharac_CPM], an algorithm to enumerate the MRs, minimal unique substrings (MUSs) [2011IlieS_MinimUniqueSubstAndMaxim] and minimal absent words (MAWs) in $O(r)$ words of compressed working space, where $r\leq n$ is the number of runs in the BWT of $T$ of length $n$.
 Given the run-length encoded BWT (RLBWT) of $T$, r-enum runs in $O(n\log\log_{w}(n/r))$ time in addition to the time linear to the number of output strings, where $w=\Theta(\log n)$ is the word size.
-We show that $O(n\log\log_{w}(n/r))$ time can be improved to $O(n)$ by resorting to the move data structure recently proposed by Nishimoto and Tabei [24] for constant-time LF-mapping on RLBWTs.
+We show that $O(n\log\log_{w}(n/r))$ time can be improved to $O(n)$ by resorting to the move data structure recently proposed by Nishimoto and Tabei [2021NishimotoT_OptimTimeQueriesOnBwt_ICALP] for constant-time LF-mapping on RLBWTs.
 We also extend r-enum to compute the NSMRs and SMRs, and the context diversity for every RMR in the same time and space complexities.
-Furthermore, we consider problems for net occurrences and net frequencies, originated from the studies of Chinese frequent strings [17, 18] and have recently attracted attention from combinatorial and algorithmic points of view [10, 28, 12, 15, 11, 21].
+Furthermore, we consider problems for net occurrences and net frequencies, originated from the studies of Chinese frequent strings [2001LinY_ExtracChinesFrequenStrinWithout, 2004LinY_ProperAndFurthApplicOf] and have recently attracted attention from combinatorial and algorithmic points of view [2024GuoEWZ_ExploitNewProperOfStrin_CPM, 2024OhlebuschBO_FasterComputOfChinesFrequen_SPIRE, 2024GuoUWZ_OnlinComputOfStrinNet_SPIRE, 2024Inenaga_FasterAndSimplOnlinComput, 2025GuoK_NetOccurInFibonAnd_CPM, 2025MienoI_SpaceEfficOnlinComputOf_CPM].
 An occurrence of a repeat is called a net occurrence if the occurrence is not covered by another repeat, and the net frequency of a repeat is the number of its net occurrences.
-We would like to point out the fact, which may have been overlooked for over two decades, that the strings with positive net frequencies are equivalent to the NSMRs [13] and the largest maximal repeats [23, 7], and thus appear to have broader applications.
-Also, net occurrences are conceptually equivalent to maximum repeats in [14].
-Given the RLBWT of $T$, we show how to compute the set $\mathcal{S}^{\mathsf{nsmr}}$ of all NSMRs in $T$ together with their net frequency/occurrences in $O(n)$ time and $O(r)$ words of space, solving the task called All-NF in [10] on RLBWTs.
-We also show that an $O(r)$-space data structure can be built from the RLBWT to support queries of computing the net frequency of any query pattern $P$ in optimal $O(|P|)$ time.
+We would like to point out the fact, which may have been overlooked for over two decades, that the strings with positive net frequencies are equivalent to the NSMRs [1997Gusfield_AlgorOnStrinTreesAnd] and the largest maximal repeats [2008NicolasRSPCDTVM_ModelLocalRepeatOnGenom, 2013Galle_BagOfRepeatRepresOf_SIGIR], and thus appear to have broader applications.
+Also, net occurrences are conceptually equivalent to maximum repeats in [2011IlieS_MinimUniqueSubstAndMaxim].
+Given the RLBWT of $T$, we show how to compute the set $\mathcal{S}^{\mathsf{nsmr}}$ of all NSMRs in $T$ together with their net frequency/occurrences in $O(n)$ time and $O(r)$ words of space, solving the task called All-NF in [2024GuoEWZ_ExploitNewProperOfStrin_CPM] on RLBWTs.
+We also address the task called Single-NF [2024GuoEWZ_ExploitNewProperOfStrin_CPM], which requires building a data structure to support NF-queries of computing the net frequency and/or net occurrences of any query pattern.
+We show that an $O(r)$-space data structure can be built from the RLBWT to support NF-queries in optimal time.
 The data structure is built in $O(r)$ space and in $O(n)$ time with high probability or deterministic $O(n+|\mathcal{S}^{\mathsf{nsmr}}|\log\log\min(\sigma,|\mathcal{S}^{\mathsf{nsmr}}|))$ time, where $\sigma\leq r$ is the alphabet size of $T$.
 To achieve this, we prove that the total number of net occurrences is less than $2r$.
-With the duality between net occurrences and MUSs shown in [14, 21], we immediately get a new upper bound $2r$ of the number of MUSs in $T$, which may be of independent interest.
+With the duality between net occurrences and MUSs shown in [2011IlieS_MinimUniqueSubstAndMaxim, 2025MienoI_SpaceEfficOnlinComputOf_CPM], we immediately get a new upper bound $2r$ of the number of MUSs in $T$, which may be of independent interest.
+1.1 Related work
+Context-sensitive repeats are closely related to and efficiently computed from suffix data structures.
+A folklore algorithm [1997Gusfield_AlgorOnStrinTreesAnd] uses suffix trees.
+To reduce the (rather big) working space of suffix trees by a constant factor, algorithms based on enhanced suffix arrays [JDA2004AbouelhodaKO_ReplacSuffixTreesWithEnhan] have been proposed.
+A compact- or compressed-space solution can be achieved by using LF-mapping and range distinct queries (see Section˜2) on BWTs or RLBWTs to simulate a traversal on suffix link trees.
+Table˜1 summarizes representative algorithms.
+reference:
+used techniques
+space (bits) running time patterns
+[1997Gusfield_AlgorOnStrinTreesAnd]: suffix tree $O(n\log n)$ $O(n\log\sigma)$ MR, NSMR, SMR
+[2009OkanoharaT_TextCategWithAllSubst_SDM]: enhanced suffix array $O(n\log n)$ $O(n)$ MR
+[2015OhlebuschB_AlphabIndepAlgorForFindin]: enhanced suffix array $O(n\log n)$ $O(n)$
+MR, NSMR, SMR,
+context diversity
+[2020BelazzouguiCKM_LinearTimeStrinIndexAnd]: BWT and RD $O(n\log\sigma)$ $O(n)$ MR, MUS, MAW
+[2012BellerBO_SpaceEfficComputOfMaxim_SPIRE]: BWT and RD $|\mathit{RD}|+O(n)$ $O(nd)$ MR, SMR
+[2015BelazzouguiC_SpaceEfficDetecOfUnusual_SPIRE]: BWT and RD $|\mathit{RD}|+O(\sigma^{2}\log^{2}n)$ $O(nd)$ MR, MUS, MAW
+[2021NishimotoT_REnumEnumerOfCharac_CPM]: RLBWT and RD $O(r\log n)$ $O(n\log\log_{w}(n/r))$ MR, MUS, MAW
+this work:
+RLBWT and RD
+$O(r\log n)$ $O(n)$
+MR, MUS, MAW,
+NSMR, SMR,
+context diversity
+Table 1: Comparison between algorithms that find characteristic substring patterns in a string $T$ of length $n$ over an integer alphabet of size $\sigma=n^{O(1)}$, where $w=\Theta(\log n)$ is the word size.
+Each method takes $T$, its BWT or $O(r)$-size RLBWT as an input, and enumerate substring patterns represented in $O(1)$ space each.
+For MAWs, we additionally need the time linear to the number of output, which is bounded by $O(n\sigma)$.
+RD represents a data structure for range distinct queries, which can be implemented in $|\mathit{RD}|$ bits of space while supporting a query in $O(d)$ time per output with $(|\mathit{RD}|,d)=(n\log_{2}\sigma+o(n\log\sigma),O(\log\sigma))$ using Lemma˜2.3 or $(|\mathit{RD}|,d)=(n\log_{2}\sigma+o(n\log\sigma),O(\log\sigma))$ [2015ClaudeNP_WavelMatrixEfficWavelTree].
+The study for net frequencies/occurrences was initiated recently by Guo et al.
+[2024GuoEWZ_ExploitNewProperOfStrin_CPM].
+Given a string $T$ of length $n$ over an integer alphabet of size $\sigma=n^{O(1)}$, we can perform All-NF in $O(n)$ time and words of space using an algorithm [2015OhlebuschB_AlphabIndepAlgorForFindin, 2024OhlebuschBO_FasterComputOfChinesFrequen_SPIRE] based on enhanced suffix arrays.
+The problem has also been studied in an online setting, where we maintain a data structure to perform All-NF and Single-NF while allowing monotonic updates to the left and/or right ends of $T$.
+The current state-of-the-art online algorithm [2025MienoI_SpaceEfficOnlinComputOf_CPM] is based on online construction of implicit suffix trees and implicit CDAWGs.
+In this paper, we achieve the first offline algorithm for All-NF and Single-NF that works in $O(r)$ words of compressed space.
+1.2 Organization of paper
+This paper is organized as follows.
+In Section˜2, we introduce notation and tools we use in the following sections.
+In Section˜3, we prove that the total number of net occurrences is less than $2r$, which is crucial to bound the time and space complexities of our algorithms for net frequencies/occurrences.
+In Section˜4, we speed up r-enum and extend it to enumerate other context-sensitive repeats.
+In Section˜5, we present an $O(r)$-space data structure for NF-queries based on RLBWTs.
+In Section˜6, we conclude the paper with some remarks.
 2 Preliminaries
@@
 We use the following known result:
-Lemma 1 ([32, 30]).
+Lemma 2.1 ([2000Willard_ExaminComputGeometVanEmde, 2008Ruzic_ConstEfficDictionInClose_ICALP]).
 Given $S\subseteq[1..u]$ of size $m\leq u\leq 2^{O(w)}$, we can build an $O(m)$-size dictionary in $O(m)$ time with high probability or deterministic $O(m\log\log m)$ time so that lookup queries can be supported in $O(1)$ worst-case time.
-A predecessor query for an integer $i$ over the set $S$ of integers asks to compute $\max\{j\in S\mid j\leq i\}$, which can be efficiently supported by the data structure of [3, Appendix A]:
-Lemma 2 ([3]).
+A predecessor query for an integer $i$ over the set $S$ of integers asks to compute $\max\{j\in S\mid j\leq i\}$, which can be efficiently supported by the data structure of [2015BelazzouguiN_OptimLowerAndUpperBound, Appendix A]:
+Lemma 2.2 ([2015BelazzouguiN_OptimLowerAndUpperBound]).
 Given $S\subseteq[1..u]$ of size $m\leq u\leq 2^{O(w)}$, we can build an $O(m)$-size data structure in $O(m\log\log_{w}(u/m))$ time and $O(m)$ space to support predecessor queries in $O(\log\log_{w}(u/m))$ time for any $i\in[1..n]$.
-A range distinct query $\mathsf{RD}(x,p,q)$ for a string $x$ and integers $1\leq p\leq q\leq|x|$ asks to enumerate $(c,p_{c},q_{c})$ such that $p_{c}$ and respectively $q_{c}$ are the smallest and largest positions in $[b..e]$ with $x[p_{c}]=x[q_{c}]=c$ for every distinct character $c$ that occurs in $x[p..q]$.
+A range distinct query $\mathsf{RD}(x,p,q)$ for a string $x$ and integers $1\leq p\leq q\leq|x|$ asks to enumerate $(c,p_{c},q_{c})$ such that $p_{c}$ and respectively $q_{c}$ are the smallest and largest positions in $[p..q]$ with $x[p_{c}]=x[q_{c}]=c$ for every distinct character $c$ that occurs in $x[p..q]$.
 It is known that a range distinct query can be supported in output-optimal time after preprocessing $x$.
-Lemma 3 ([22, 2]).
+Lemma 2.3 ([2002Muthukrishnan_EfficAlgorForDocumRetriev_SODA, 2020BelazzouguiCKM_LinearTimeStrinIndexAnd]).
 Given a string $x\in[1..\sigma]^{m}$, we can build an $O(m)$-size data structure in $O(m)$ time and space to support range distinct queries.
 With a null-initialized array of size $\sigma$, each query can be answered in $O(k)$ time, where $k$ is the number of outputs.
-A minor remark is that the output of range distinct queries of Lemma˜3 may not be sorted in any order of characters.
+A minor remark is that the output of range distinct queries of Lemma˜2.3 may not be sorted in any order of characters.
 2.3 Context-sensitive repeats
@@
 We set the following exception for the definition of net occurrences of $\varepsilon$: An occurrence $b\in\mathsf{Occ}(\varepsilon)$ is a net occurrence of $\varepsilon$ if and only if T[b1]T[b-1] and $T[b]$ are both unique.
-This exception provides a smooth transition to another characterization of net occurrences based on the uniqueness of extended net occurrences [10, 21].
+This exception provides a smooth transition to another characterization of net occurrences based on the uniqueness of extended net occurrences [2024GuoEWZ_ExploitNewProperOfStrin_CPM, 2025MienoI_SpaceEfficOnlinComputOf_CPM].
 Table 3 summarizes the concepts introduced in this subsection for $T=\mathtt{abcbbcbcabc\mathdollar}$.
@@
 A link from a node $x$ to a (possibly implicit) node $ax\in\mathcal{S}$ for some $a\in\Sigma_{\mathdollar}$ is called a Weiner link.
-It is known that every node can be reached by following Weiner links from the root $\varepsilon$, and the total number of Weiner links is at most $3n$, which can be seen from the duality between suffix trees and DAWGs [5].
+It is known that every node can be reached by following Weiner links from the root $\varepsilon$, and the total number of Weiner links is at most $3n$, which can be seen from the duality between suffix trees and DAWGs [1985BlumerBHECS_SmallAutomRecogSubworOf].
 The suffix array $\mathsf{SA}[1..n]$ of $T$ is the integer array such that, for any $i\in[1..n]$, $T[\mathsf{SA}[i]..]$ is the lexicographically $i$-th suffix among the non-empty suffixes of $T$.
@@
 The Burrows-Wheeler transform (BWT) $\mathsf{L}[1..n]$ of $T$ is the string such that $\mathsf{L}[i]=T[\mathsf{SA}[i]-1]$ for any $i\in[1..n]$.
+Note that $T[\mathsf{SA}[i]-1]=T[0]=\mathdollar$ for $i$ with $\mathsf{SA}[i]=1$ due to the sentinel at $T[0]$.
 By definition, it holds that $\mathsf{lc}(x)=\{\mathsf{L}[i]\mid i\in\mathcal{I}(x)\}$ for any string $x$.
@@
 The FL-mapping $\mathsf{FL}$ is the inverse mapping of $\mathsf{LF}$.
-Given an integer $i$ in $\mathcal{I}(x)$ for a string $x$, we can compute $x$ by using FL-mapping $|x|$ times from $i$ because $x[k]=\mathsf{L}[\mathsf{FL}^{k}(i)]$ holds for any $k~(1\leq k\leq|x|)$, where $\mathsf{FL}^{k}(i)$ is defined recursively $\mathsf{FL}^{k}(i)=\mathsf{FL}(\mathsf{FL}^{k-1}(i))$ with the base case $\mathsf{FL}^{1}(i)=\mathsf{FL}(i)$.
-Table˜2 shows an example of $\mathsf{SA}$, $\mathsf{LF}(\cdot)$, $\mathsf{FL}(\cdot)$ and $\mathsf{L}$ for $T=\mathtt{abcbbcbcabc\mathdollar}$.
+Given an integer $i$ in $\mathcal{I}(x)$ for a string $x$, we can compute $x$ by using FL-mapping $|x|$ times from $i$ because $x[k]=\mathsf{L}[\mathsf{FL}^{k}(i)]$ holds for any $k\penalty 10000\ (1\leq k\leq|x|)$, where $\mathsf{FL}^{k}(i)$ is defined recursively $\mathsf{FL}^{k}(i)=\mathsf{FL}(\mathsf{FL}^{k-1}(i))$ with the base case $\mathsf{FL}^{1}(i)=\mathsf{FL}(i)$.
+Table˜3 shows an example of $\mathsf{SA}$, $\mathsf{LF}(\cdot)$, $\mathsf{FL}(\cdot)$ and $\mathsf{L}$ for $T=\mathtt{abcbbcbcabc\mathdollar}$.
 Figure˜1 also illustrates the suffix tree of $T=\mathtt{abcbbcbcabc\mathdollar}$ and a relationship between Weiner links and BWTs.
@@
 While $r\leq n$ is obvious, inequality $\sigma\leq r$ holds under the assumption that $T$ contains all characters in $\Sigma_{\mathdollar}$.
-Example 4.
+Example 2.4.
 For $T=\mathtt{abcbbcbcabc\mathdollar}$, its BWT $\mathsf{L}=\mathtt{cc\mathdollar cacabbbbb}$ has $r=7$ runs.
@@
 Hence, LF-mapping can be implemented in $O(r)$ space by constructing the predecessor data structure for the set of beginning positions of runs of $\mathsf{L}$ and remembering the $\mathsf{LF}(\cdot)$ values for those positions (FL-mapping can be implemented similarly).
-Using the predecessor data structure of Lemma˜2, we get:
-Lemma 5.
+Using the predecessor data structure of Lemma˜2.2, we get:
+Lemma 2.5.
 Given the RLBWT of $T$, we can construct an $O(r)$-size data structure in $O(r\log\log_{w}(n/r))$ time and $O(r)$ space to compute $\mathsf{LF}(i)$ and $\mathsf{FL}(i)$ in $O(\log\log_{w}(n/r))$ time for any $i\in[1..n]$.
-Nishimoto and Tabei [24] proposed a novel data structure, called the move data structure, to implement LF-mapping by a simple linear search on a list of intervals without predecessor queries.
+Nishimoto and Tabei [2021NishimotoT_OptimTimeQueriesOnBwt_ICALP] proposed a novel data structure, called the move data structure, to implement LF-mapping by a simple linear search on a list of intervals without predecessor queries.
 We consider a list $p_{1}..p_{2}),[p_{2}..p_{3}),\dots,[p_{\hat{r}}..p_{\hat{r}+1})$ of $\hat{r}=O(r)$ intervals such that $p_{1}=1$, $p_{\hat{r}+1}=n+1$, and for any $k\in[1..\hat{r}$], $\mathsf{L}[p_{k}..p_{k+1})$ consists of a single character and $\mathsf{LF}(p_{k})..\mathsf{LF}(p_{k+1}-1)]$ intersects with $O(1)$ intervals in the list.
@@
 Note that $\mathsf{LF}(i)$ computation on the move data structure works along with the pointer to the interval in the list that contains $i$, which is handled implicitly hereafter.
-Under the assumption that $\sigma\leq r$ in this paper, we can construct the data structure needed for LF-mapping (and FL-mapping) in $O(r\log r)$ time [25, Appendix D]:
-Lemma 6.
+Under the assumption that $\sigma\leq r$ in this paper, we can construct the data structure needed for LF-mapping (and FL-mapping) in $O(r\log r)$ time [2020NishimotoT_FasterQueriesOnBwtRuns_X3, Appendix D]:
+Lemma 2.6.
 Given the RLBWT of $T$, we can construct an $O(r)$-size data structure in $O(r\log r)$ time and $O(r)$ space to compute $\mathsf{LF}(i)$ and $\mathsf{FL}(i)$ in $O(1)$ time for any $i\in[1..n]$.
@@
 We start with the following lemma.
-Lemma 7.
+Lemma 3.1.
 For any integer $i\in[p..q]=\mathcal{I}(x)$ for a repeat $x$ of $T$, $\mathsf{SA}[i]$ is a net occurrence of $x$ if and only if $\mathsf{L}[i]$ is unique in $\mathsf{L}[p..q]$ and $T[\mathsf{SA}[i]..\mathsf{SA}[i]+|x|]$ is unique in $T$.
-Proof.
+Proof 3.2.
 If $a=\mathsf{L}[i]$ is not unique in $\mathsf{L}[p..q]$, then $\mathsf{SA}[i]$ is not a net occurrence of $x$ because $ax=T[\mathsf{SA}[i]-1..\mathsf{SA}[i]+|x|)$ is a repeat that covers $\mathsf{SA}[i]..\mathsf{SA}[i]+|x|)$.
@@
 On the other hand, if $\mathsf{L}[i]$ is unique in $\mathsf{L}[p..q]$ and $T[\mathsf{SA}[i]..\mathsf{SA}[i]+|x|]$ is unique in $T$, there is no repeat that covers $\mathsf{SA}[i]..\mathsf{SA}[i]+|x|)$, and hence, $\mathsf{SA}[i]$ is a net occurrence of $x$.
-∎
-Example 8.
-Let us consider a repeat $\mathtt{bc}$ with $\mathcal{I}(\mathtt{bc})=[5..8]$ in the example of Table˜2.
+Example 3.3.
+Let us consider a repeat $\mathtt{bc}$ with $\mathcal{I}(\mathtt{bc})=[5..8]$ in the example of Table˜3.
 $\mathsf{SA}[6]=7$ is a net occurrence of $\mathtt{bc}$ because $\mathsf{L}[6]=\mathtt{c}$ is unique in $\mathsf{L}[5..8]=\mathtt{acab}$, and $T[\mathsf{SA}[6]..\mathsf{SA}[6]+2]=T[7..9]=\mathtt{bca}$ is unique in $T$.
@@
 $\mathsf{SA}[7]=2$ and $\mathsf{SA}[8]=5$ are not net occurrences of $\mathtt{bc}$ because $T[2..4]=T[5..7]=\mathtt{bcb}$ is a repeat.
-The following lemma was implicitly proved in [10, Theorem 24] to show the total length of all net occurrences is bounded by twice the sum of irreducible LCPs.
-Lemma 9.
+The following lemma was implicitly proved in [2024GuoEWZ_ExploitNewProperOfStrin_CPM, Theorem 24] to show the total length of all net occurrences is bounded by twice the sum of irreducible longest common prefixes.
+Lemma 3.4.
 If $\mathsf{SA}[i]$ is a net occurrence of a repeat $x$ of $T$, then $i$ is at the beginning or ending position of a run in $\mathsf{L}$, and $\mathsf{SA}[i]$ cannot be a net occurrence of another repeat.
-Proof.
+Proof 3.5.
 Let $p..q]=\mathcal{I}(x)$.
-If $i\in[p..q]$ is not at the run’s boundaries of $\mathsf{L}$, then at least one of $\mathsf{L}[i]=\mathsf{L}[i+1]$ with $i+1\in[p..q]$ or $\mathsf{L}[i-1]=\mathsf{L}[i]$ with $i-1\in[p..q]$ holds, which means that $\mathsf{SA}[i]$ is not a net occurrence of $x$ by Lemma˜7.
+If $i\in[p..q]$ is not at the run’s boundaries of $\mathsf{L}$, then at least one of $\mathsf{L}[i]=\mathsf{L}[i+1]$ with $i+1\in[p..q]$ or $\mathsf{L}[i-1]=\mathsf{L}[i]$ with $i-1\in[p..q]$ holds, which means that $\mathsf{SA}[i]$ is not a net occurrence of $x$ by Lemma˜3.1.
 $\mathsf{SA}[i]$ cannot be a net occurrence of two distinct repeats, since otherwise, the occurrence of the shorter repeat is covered by the longer one.
-∎
-Lemma˜9 leads to the following theorem:
-Theorem 10.
+Lemma˜3.4 leads to the following theorem:
+Theorem 3.6.
 The total number of net occurrences of $T$ is less than $2r$.
-Proof.
+Proof 3.7.
 Let $r_{1}$ and respectively $r_{>1}$ be the number of runs in $\mathsf{L}$ of length $1$ and longer than $1$, i.e., $r=r_{1}+r_{>1}$.
-Since there are $r_{1}+2r_{>1}$ positions that correspond to beginning or ending positions of runs in $\mathsf{L}$, it is clear from Lemma˜9 that the total number of net occurrences is at most $r_{1}+2r_{>1}\leq 2r$.
+Since there are $r_{1}+2r_{>1}$ positions that correspond to beginning or ending positions of runs in $\mathsf{L}$, it is clear from Lemma˜3.4 that the total number of net occurrences is at most $r_{1}+2r_{>1}\leq 2r$.
 The number is shown to be less than $2r$ by looking closer at some special positions like $1$ and $n$ on $\mathsf{L}$: Since $\mathsf{SA}[1]$ (resp.
@@
 last) run of $\mathsf{L}$.
-∎
 A non-empty interval $[b..e]$ is called a minimal unique substring (MUS) if $T[b..e]$ is unique while $T[b+1..e]$ and T[b..e1]T[b..e-1] are repeats.
-In [14, 21], it has been proved that the net occurrences and MUSs are dual concepts, basically saying that for any two consecutive MUSs $[b..e]$ and $b^{\prime}..e^{\prime}]$ with $b<b^{\prime}$, $b+1$ is a net occurrence of $T[b+1..e^{\prime}-1]$.
+In [2011IlieS_MinimUniqueSubstAndMaxim, 2025MienoI_SpaceEfficOnlinComputOf_CPM], it has been proved that the net occurrences and MUSs are dual concepts, basically saying that for any two consecutive MUSs $[b..e]$ and $b^{\prime}..e^{\prime}]$ with $b<b^{\prime}$, $b+1$ is a net occurrence of $T[b+1..e^{\prime}-1]$.
 With this duality, we immediately get the following:
-Corollary 11.
+Corollary 3.8.
 The number of MUSs of $T$ is less than $2r$.
@@
 Each node $x\in\mathcal{S}^{\mathsf{rmr}}$ is visited with its suffix interval $\mathcal{I}(x)$ and length $|x|$ together with the information about its right extensions $\mathsf{rlist}(x)$, which is the list of the set $\{(c,\mathcal{I}(xc))\mid c\in\mathsf{rc}(x)\}$ sorted by the character $c$.
-The triplet $\mathsf{repr}(x)=(\mathcal{I}(x),\mathsf{rlist}(x),|x|)$ is called the rich representation [2, 26] of $x$, which can be stored in $O(|\mathsf{rc}(x)|)$ space.
+The triplet $\mathsf{repr}(x)=(\mathcal{I}(x),\mathsf{rlist}(x),|x|)$ is called the rich representation [2020BelazzouguiCKM_LinearTimeStrinIndexAnd, 2021NishimotoT_REnumEnumerOfCharac_CPM] of $x$, which can be stored in $O(|\mathsf{rc}(x)|)$ space.
 Given $\mathsf{repr}(x)=([p..q],\{(c_{i},[p_{i}..q_{i}])\}_{i=1}^{k},|x|)$, we can compute $\{\mathsf{repr}(ax)\mid ax\in\mathcal{S},a\in\Sigma\}$ as follows: A null-initialized array $A[1..\sigma]$ is used as a working space to build $\mathsf{rlist}(ax)$ at $A[a]$ for $a\in\mathsf{lc}(x)\setminus\{\mathdollar\}$.
@@
 For the next round of the breadth-first traversal on $\mathcal{S}^{\mathsf{rmr}}$, we keep only the set $\{\mathsf{repr}(ax)\mid ax\in\mathcal{S}^{\mathsf{rmr}}\}$, discarding $\mathsf{repr}(ax)$ if $\mathsf{rlist}(ax)$ contains only one element.
-Example 12.
+Example 4.1.
 In the r-enum for $T=\mathtt{abcbbcbcabc\mathdollar}$, we visit a repeat $\mathtt{bc}\in\mathcal{S}^{\mathsf{rmr}}$ with its rich representation $\mathsf{repr}(\mathtt{bc})=(\mathcal{I}(\mathtt{bc}),\mathsf{rlist}(\mathtt{bc}),|\mathtt{bc}|)=([5..8],\{(\mathdollar,[5..5]),(\mathtt{a},[6..6]),(\mathtt{b},[7..8])\},2)$.
@@
 R-enum executes the above procedure in $O(r)$ space using range distinct queries on $\mathsf{L}^{\prime}[1..r]$ and RLBWT-based LF-mapping.
-The breadth-first traversal on $\mathcal{S}^{\mathsf{rmr}}$ can be made in $O(r)$ space because the space needed to store the rich representations for all strings in $\{ax\in\mathcal{S}\mid a\in\Sigma,x\in\mathcal{S}^{\mathsf{rmr}},|x|=t\}$ for any $t\in[0..n]$ was proved to be $O(r)$ in Section 3.3 of [26].
+The breadth-first traversal on $\mathcal{S}^{\mathsf{rmr}}$ can be made in $O(r)$ space because the space needed to store the rich representations for all strings in $\{ax\in\mathcal{S}\mid a\in\Sigma,x\in\mathcal{S}^{\mathsf{rmr}},|x|=t\}$ for any $t\in[0..n]$ was proved to be $O(r)$ in Section 3.3 of [2021NishimotoT_REnumEnumerOfCharac_CPM].
 The total number of outputs of range distinct queries is upper bounded by $O(n)$ because each output can be charged to a distinct Weiner link.
 Similarly, the total number of LF-mapping needed is bounded by $O(n)$.
-The original r-enum used Lemma˜5 to implement LF-mapping and runs in $O(n\log\log_{w}(n/r))$ time in total (excluding the time to output patterns for MAWs).
+The original r-enum used Lemma˜2.5 to implement LF-mapping and runs in $O(n\log\log_{w}(n/r))$ time in total (excluding the time to output patterns for MAWs).
 We show that this can be improved to $O(n)$ time.
-Theorem 13.
+Theorem 4.2.
 Given the RLBWT of $T$, r-enum runs in $O(r)$ working space and $O(n)$ time in addition to the time linear to the number of output strings.
-Proof.
-If $r=\Omega(n/\log n)$, the original r-enum with Lemma˜5 is already upper bounded by $O(n)$ because $O(n\log\log_{w}(n/r))=O(n\log\log_{w}\log n)$ and $w=\Theta(\log n)$.
-For the case with $r=o(n/\log n)$, we use the move data structure of Lemma˜6 to implement LF-mapping, which can be constructed in $O(r\log r)=O(n)$ time.
+Proof 4.3.
+If $r=\Omega(n/\log n)$, the original r-enum with Lemma˜2.5 is already upper bounded by $O(n)$ because $O(n\log\log_{w}(n/r))=O(n\log\log_{w}\log n)$ and $w=\Theta(\log n)$.
+For the case with $r=o(n/\log n)$, we use the move data structure of Lemma˜2.6 to implement LF-mapping, which can be constructed in $O(r\log r)=O(n)$ time.
 Since LF-mapping can be done in $O(1)$ time on the move data structure, the r-enum runs in $O(n)$ time.
-∎
 4.2 Extension for NSMRs, SMRs, and context diversities
 In this subsection, we extend r-enum to compute the NSMRs and SMRs, and the context diversity for every RMR.
+Since $\mathcal{S}^{\mathsf{smr}}\subseteq\mathcal{S}^{\mathsf{nsmr}}\subseteq\mathcal{S}^{\mathsf{mr}}\subseteq\mathcal{S}^{\mathsf{rmr}}$ holds, for enumerating SMRs (resp.
+NSMRs), it is enough to check every element in $\mathcal{S}^{\mathsf{rmr}}$ if it is in $\mathcal{S}^{\mathsf{smr}}$ (resp.
+$\mathcal{S}^{\mathsf{nsmr}}$).
 Recall that r-enum traverses $x\in\mathcal{S}^{\mathsf{rmr}}$ while computing $\{\mathsf{repr}(ax)\mid ax\in\mathcal{S}\}$ from $\mathsf{repr}(x)=([p..q],\{(c_{i},[p_{i}..q_{i}])\}_{i=1}^{k},|x|)$.
@@
 Since we can check the conditions for NSMRs and SMRs along with the traversal of $\mathcal{S}^{\mathsf{rmr}}$, the time and space complexities of r-enum are retained.
-Theorem 14.
+Theorem 4.4.
 Given the RLBWT of $T$, we can compute the NSMRs and SMRs, and the context diversity for every RMR in $O(n)$ time and $O(r)$ working space.
@@
 5 An $O(r)$-size data structure for NF-queries
-In this section, we tackle the problem called Single-NF [10] on RLBWTs: We consider preprocessing the RLBWT of $T$ to construct a data structure of size $O(r)$ so that we can support NF-queries of computing $\mathsf{NF}(P)$ for any query pattern $P$.
+In this section, we tackle the problem called Single-NF [2024GuoEWZ_ExploitNewProperOfStrin_CPM] on RLBWTs: We consider preprocessing the RLBWT of $T$ to construct a data structure of size $O(r)$ so that we can support NF-queries of computing $\mathsf{NF}(P)$ for any query pattern $P$.
 We define the reversed trie for a set $S$ of strings as follows:
@@
   • There is an edge from node $x$ to $ax$ with a character $a$.
-Theorem 15.
+Theorem 5.1.
 Given the RLBWT of $T$, we can build an $O(r)$-size data structure in $O(r)$ space and in $O(n)$ time with high probability or deterministic $O(n+|\mathcal{S}^{\mathsf{nsmr}}|\log\log\min(\sigma,|\mathcal{S}^{\mathsf{nsmr}}|))$ time to support NF-queries for any query pattern $P$ in $O(|P|)$ time.
-Proof.
-We execute r-enum of Theorem˜14, which actually performs a breadth-first traversal of the reversed trie for $\mathcal{S}^{\mathsf{rmr}}$ in $O(n)$ time and $O(r)$ space.
-Our idea is to build the compacted reversed trie $\mathcal{T}$ for $\mathcal{S}^{\mathsf{nsmr}}$ during the traversal, which can be stored in $O(|\mathcal{S}^{\mathsf{nsmr}}|)=O(r)$ space because $|\mathcal{S}^{\mathsf{nsmr}}|\leq 2r$ by Lemma˜9.
+Proof 5.2.
+We execute r-enum of Theorem˜4.4, which actually performs a breadth-first traversal of the reversed trie for $\mathcal{S}^{\mathsf{rmr}}$ in $O(n)$ time and $O(r)$ space.
+Our idea is to build the compacted reversed trie $\mathcal{T}$ for $\mathcal{S}^{\mathsf{nsmr}}$ during the traversal, which can be stored in $O(|\mathcal{S}^{\mathsf{nsmr}}|)=O(r)$ space because $|\mathcal{S}^{\mathsf{nsmr}}|\leq 2r$ by Lemma˜3.4.
 For each node $x$ of $\mathcal{T}$, which corresponds to a string in $\mathcal{S}^{\mathsf{nsmr}}$ or a branching node, we store the string length $|x|$ and an integer $i_{x}$ in $\mathcal{I}(x)$ so that the edge label between $x$ and its parent $y$ can be retrieved by using FL-mapping |x||y||x|-|y| times from $i_{x}$.
@@
 We also remove $y$, if it becomes a non-branching node not in $\mathcal{S}^{\mathsf{nsmr}}$.
-In total, these tasks can be done in $O(n)$ time and $O(r)$ space along with r-enum of Theorem˜14.
-For each branching node of $\mathcal{T}$ with $k\leq\min(\sigma,|\mathcal{S}^{\mathsf{nsmr}}|)$ outgoing edges, we build a dictionary of Lemma˜1 in $O(k)$ space and $O(k)$ time w.h.p.
+In total, these tasks can be done in $O(n)$ time and $O(r)$ space along with r-enum of Theorem˜4.4.
+For each branching node of $\mathcal{T}$ with $k\leq\min(\sigma,|\mathcal{S}^{\mathsf{nsmr}}|)$ outgoing edges, we build a dictionary of Lemma˜2.1 in $O(k)$ space and $O(k)$ time w.h.p.
 or deterministic $O(k\log\log\min(\sigma,k))$ time so that we can decide which child to proceed by the first characters of edges in $O(1)$ time.
@@
 If we can reach a node $P\in\mathcal{S}^{\mathsf{nsmr}}$, we output $\mathsf{NF}(P)$ stored in the node.
-∎
-We can also support queries of computing the net occurrences of a query pattern $P$ in $O(|P|+|\mathsf{NOcc}(P)|)$ time by storing $\mathsf{NOcc}(x)$ for each node $x\in\mathcal{S}^{\mathsf{nsmr}}$ in $\mathcal{T}$, which can be stored in $O(r)$ space in total due to Lemma˜9.
+We can also support queries of computing the net occurrences of a query pattern $P$ in $O(|P|+|\mathsf{NOcc}(P)|)$ time by storing $\mathsf{NOcc}(x)$ for each node $x\in\mathcal{S}^{\mathsf{nsmr}}$ in $\mathcal{T}$, which can be stored in $O(r)$ space in total due to Lemma˜3.4.
 Figure 2: An illustration of the compacted reversed trie $\mathcal{T}$ for $\mathcal{S}^{\mathsf{nsmr}}=\{\mathtt{bc},\mathtt{abc},\mathtt{bcb}\}$ in our running example $T=\mathtt{abcbbcbcabc\mathdollar}$.
@@
 We improved the time complexity of r-enum, and extended its enumeration power for other context-sensitive repeats in the literature.
-We also proposed the first data structure of size $O(r)$ to support NF-queries in optimal time.
+We also proposed the first data structure of size $O(r)$ to support NF-queries in optimal time, which can be regarded as an $O(r)$-space representation of $\mathcal{S}^{\mathsf{nsmr}}$ with the help of RLBWTs.
 As a combinatorial property, we formally proved that the total number of net occurrences, as well as MUSs, is upper bounded by $O(r)$.
-We finally remark that our work leads to the first algorithm to compute the sorted list of net occurrences and MUSs in $O(n)$ time and $O(r)$ space: We enumerate all net occurrences by Theorem˜13, and sort them in $O(n)$ time and $O(r)$ space by [25, Lemma 13].
-Acknowledgements
-Tomohiro I was supported by JSPS KAKENHI (Grant Numbers JP24K02899) and JST AIP Acceleration Research JPMJCR24U4, Japan.
+We note that our enumeration algorithm runs independently of the alphabet size $\sigma$ thanks to the output-optimal range distinct queries of Lemma˜2.3.
+The space usage depending on $\sigma$ is hidden behind $O(r)$.
+We removed the dependence on $\sigma$ in the NF-query time by using a dictionary with worst-case constant-time lookup.
+Finally, we remark that our work leads to the first algorithm to compute the sorted list of net occurrences and MUSs in $O(n)$ time and $O(r)$ space: We enumerate all net occurrences by Theorem˜4.2, and sort them in $O(n)$ time and $O(r)$ space by [2020NishimotoT_FasterQueriesOnBwtRuns_X3, Lemma 13].